Planning week 9-16 February 2013

This was a quiet week, with little flaring activity and no HOPs. The default pointing was AR 11670, a small AR with two sunspots that produced occasional C-flares. It was just to the north of a large, decaying active region 11669 which had nice fan loops.

For OP3 I requested that the AR pointing be changed to the new AR coming around the east limb. It had recently emerged on 11 February so I was hoping that it may show some activity.

Saturday 9 February (OP 1)

AIA 193 image

There was a nice coronal cavity on the SE limb, so HOP 114 was scheduled from 10:29-13:00. The EIS study was GDZ_PLUME1_2_300_50s. The pointing then shifted to the west limb where I ran a test of HPW023 at 13:05.

The default target for the rest of the OP was then AR 11670. For 15:00-18:14 I ran a test of PRY_flare_1 (a new flare study), with little hope of catching a flare.

From 21:37 I ran PRY_loop_velocity for a large fan loop system.

Sunday 10 February (OP1)

AIA 193 image

The fan loop study continued until 02:13, and then I ran Atlas_60 at the base of the fan loops at 02:17.

As part of general AR monitoring, I ran HPW021 over the core of AR 11669 (large decaying AR) at 06:57.

I ran PRY_flare_1 on AR 11670 again at 09:11 for 8.5 hours, but the Sun was very quiet in this timefram. I then returned to AR 11669 for another HPW021 at 22:14.

Monday 11 February (OP 1)

AIA 193 image

To pad out the plan I scheduled the Flare266_Hunter01 study at 07:40 (5 hours) and again at 13:00 (4 hours). I set the response to be HPW009 (sit-and-stare). There was a B7 flare at about 16:00, but I doubt this triggered anything.

Tuesday 12 February (OP 2)

AIA 193 image

I just scheduled Cam_flare_diag (small flare study) for 10:13-02:17.

Wednesday 13 February (OP 2)

AIA 193 image

AR 11634 (southern hemisphere) returned and so I scheduled HPW021 at 06:15. It turns out some new flux emerged within the old AR (which is very diffuse), and so the target was actually AR 11672.

Pointing then switched to quiet Sun near disk center for Atlas_120 and SYNOP006 for 09:00-15:00.

Thursday 14 February (OP 3)

AIA 193 image

A spectral atlas was run at the east limb at 09:49. There may be some contamination from weak AR loops.

At 11:02 I began a very long run of Tongjiang's flare oscillation study (hot_loop_osc), running until 03:20 on 15 Feb. The target was an AR that just appeared around the east limb. The loops seem quite dynamic in AIA images, but I'm sceptical that there will be any flaring.

Friday 15 February (OP 3)

AIA 193 image

HPW021 (part of the regular NRL AR monitoring sequence) was run on AR 11671 at 03:38.

I finished the plan by scheduling Tongjiang's study again from 07:42 for about 9 hours. The target is the off-limb AR again.

Saturday 16 February (OP 3)

No science studies performed.